OpenStudy (anonymous):

find the Maclaurin series for (x^2)cos(x). find the interval of convergence of the series.

OpenStudy (anonymous):

\[\cos x = \sum_{n=0}^{\infty} (-1)^n {{x^{2n}} \over {(2n)!}}\] so\[ x^2\cos x = \sum_{n=0}^{\infty} (-1)^n {{x^{2n+2}} \over {(2n)!}}\]

OpenStudy (anonymous):

interval of convergence \[(-\infty,\infty)\]